OpenStudy (starfireandrobin):

Help!! How do you solve 2m-4/2??

OpenStudy (jackthegreatest):

\[\frac{ 2m-4 }{ 2 }\]

OpenStudy (jackthegreatest):

is that right?

OpenStudy (starfireandrobin):

\[\frac{ 2m-4 }{ 2 }\]

OpenStudy (starfireandrobin):

Ah, yes lol

OpenStudy (jackthegreatest):

does it equal anything?

OpenStudy (starfireandrobin):

Well, the whole equation is 6m +3 = 2m-4/2, but I'm just trying to solve this right side right now.

OpenStudy (jackthegreatest):

lol should have said that first

OpenStudy (starfireandrobin):

To then hopefully solve the whole thing. Because I think I'll get it afterwards. I just confused with this.

OpenStudy (jackthegreatest):

well it simplifies to m-2

OpenStudy (starfireandrobin):

I put it through a calculator, and it came out 6m + 3 = m -2, but I don't know why

OpenStudy (starfireandrobin):

I just need to understand how I get m - 2.

OpenStudy (jackthegreatest):

2(m-2)/2

OpenStudy (jackthegreatest):

take out the 2s and you get m-2

OpenStudy (jackthegreatest):

do you know how to factor?

OpenStudy (starfireandrobin):

I'm still confused. And I'm not sure if I do.

OpenStudy (starfireandrobin):

What happens to the 4??

OpenStudy (jackthegreatest):

2 x 2 is 4 right

OpenStudy (starfireandrobin):

Yep.

OpenStudy (jackthegreatest):

m times 2 is 2m right

OpenStudy (starfireandrobin):

Yep.

OpenStudy (jackthegreatest):

so 2m-4 = 2(m-2)

OpenStudy (starfireandrobin):

So we're just reversing what was done???

OpenStudy (jackthegreatest):

im not sure what you mean

OpenStudy (jackthegreatest):